# appendices

### v1

@@ 5 Evaluation
To comprehensively study the performance of CTINexus
in various phases of CSKG construction, we set the following four research questions:
@@ 1. RQ1:
    How does CTINexus’s performance in extracting cybersecurity entities and relations compare to existing baseline methods?
@@ 2. RQ2:
    How well does CTINexus perform in cybersecurity triplet extraction?
@@ 3. RQ3:
    How well does CTINexus perform in knowledge graph construction?
@@ 4. RQ4:
5.1 Dataset Annotation
Existing datasets benchmark triplet extraction but do not encompass other procedures in our pipeline, and the CTI reports they include are mostly outdated. For example, the dataset constructed by LADDER is limited to reports from 2010 to 2021 [30]. To address this, we created a dataset to evaluate CTINexus across cybersecurity triplet extraction, hierarchical entity alignment, and long-distance relation prediction phases. Our dataset includes 150 reports from May 2023 onwards, Published by organizations like Trend Micro, Symantec, and The Hacker News, it includes 10 sources, averaging 15 reports each. Annotators with over three years of threat analysis expertise followed a four-phase process: Phase I involved annotating all cybersecurity entities and selecting entity types; Phase II identified explicit relations among entities and organized them into JSON-formatted triplets; Phase III grouped entities by type and merged those referring to the same threat-related concept; Phase IV selected central and topic entities based on degree and summarized implicit relations. This process resulted in 59,776 mentions, 35,258 entities, and 34,876 relations, enabling an effective evaluation of CTINexus ’s performance in constructing cohesive and complete CSKGs.
@@ 5.1 Dataset Annotation
5.2 RQ1: How does CTINexus compare against existing CTI extraction methods?
@@
We evaluate CTINexus against two state-of-the-art baselines: EXTRACTOR [71] and LADDER [30], representing syntactic analysis-based and fine-tuning-based approaches, respectively. Several methodological challenges were addressed to enable fair comparison. For EXTRACTOR, we adapted its output to our broader ontology using CTINexus’s coarse-grained entity grouping module. For LADDER, we addressed two key differences: (1) LADDER uses a word-level annotation format, where each token is labeled with its target class. In contrast, our dataset follows an end-to-end report-to-triplet format, where the entire report is input, and the label is a set of extracted triplets. (2) LADDER uses a simplified ontology derived from MALONT, which includes only 10 entity types, a subset of the entity types used in our ontology. To facilitate comparison, we developed scripts to tokenize our data and convert our manually annotated datasets into LADDER’s word-level format. To ensure a fair comparison with LADDER, we merged our training set with LADDER’s in a 5:1 ratio, maintaining their original training/validation split. We also replaced LADDER’s test set with ours to ensure consistent evaluation on the same data. We compare with LADDER solely on named entity extraction performance. The reason is that our method focuses on open relation extraction, while LADDER targets relation classification within fixed categories.
@@
Table I demonstrates that CTINexus outperforms EXTRACTOR in all metrics in cybersecurity triplet extraction. The evaluation results in Section 5.3 showed that GPT-4 outperforms all other backbone models. Thus, we use GPT-4 as the default backbone model for CTINexus’s implementation. This superior performance can be attributed to several factors. First, CTINexus leverages the robust context understanding and instruction-following capabilities of LLMs and enhances specificity with kNN-selected demonstration examples for extracting triplets. In contrast, EXTRACTOR employs general fine-tuning to extract semantic roles not specific to any ontology, reducing its accuracy in triplet extraction. Also, the CTI context introduces peculiarities that lead to errors in EXTRACTOR’s semantic role labeling module, which relies on a simple BERT model. For instance, EXTRACTOR might extract a triplet like $\langle$“Androxgh0st malware”, “support”, “numerous functions capable of abusing the Simple Mail Transfer Protocol (SMTP), such as scanning and exploiting exposed credentials and application programming interfaces (APIs), and web shell deployment”$\rangle$, where the object is a long sentence not suitable as a single entity. The object contains multiple entities due to misidentified boundaries. Conversely, CTINexus can comprehend implicit meanings and transform phrases to be more suitable as entities, resulting in a triplet like $\langle$“Androxgh0st malware”, “supports”, “functions abusing SMTP”$\rangle$.
@@
Table II demonstrates that CTINexus outperforms LADDER in F1 Score, Precision, and Recall by 26.7%, 17.5%, and 19.5%, respectively. Specifically, LADDER achieved an F1 Score of 71.13%, Precision of 78.31%, and Recall of 73.94%, which are slightly lower than the numbers reported in LADDER’s original evaluation (75.32%, 79.06%, and 76.98%, respectively). LADDER’s lower performance on our test data compared to its reported values is likely due to a distribution shift. LADDER’s dataset spans 2015 to 2021, while our data is from May 2023 onward. This temporal gap may introduce new patterns, terminologies, or threat vectors that LADDER’s model struggles to generalize to, even when retrained on a mix of old and new data. The performance disparity between LADDER and CTINexus can be attributed to several factors. First, fine-tuning the model in LADDER may lead to overfitting on the training set. Consequently, when confronted with unseen entities in the test set, the model may struggle to recognize them accurately, potentially misclassifying them or recognizing only parts of the entities. For example, in the sentence “… with a specific focus on WordPress sites”, LADDER extracts only “WordPress” as an application, resulting in ambiguous content. In contrast, CTINexus correctly extracts “WordPress sites”, which more accurately reflects the original context. Second, similar to EXTRACTOR, the LADDER model lacks sufficient contextual understanding. For instance, in the sentence “The victims include Family Day Care Services, a Canadian childcare service”, LADDER incorrectly identifies “Canadian” as a “B-Location”, whereas it should be recognized as a descriptive term for the childcare service. Furthermore, LADDER models relation extraction as relation classification, limited to ten relation classes (i.e., closed-world setting). This constraint restricts the contextual information in the extracted content and hinders the model’s ability to generalize to new CTI data containing different or additional relation classes.
@@
Method F1 Score Precision Recall
@@
Method F1 Score Precision Recall
@@
5.3 RQ2: How well does CTINexus perform in cybersecurity triplet extraction
To demonstrate the effectiveness of CTINexus in cybersecurity triplet extraction, stemming from the superiority of the ICL paradigm and our specific prompt design, we conducted experiments on different ICL configurations, focusing on three aspects: (1) the number of demonstration examples, (2) the permutation of these examples, and (3) the backbone model types. By default, CTINexus uses GPT-4 as the model backbone, selects the $k$ most similar prompt examples sorting in ascending order of query similarity.
@@ 5.3 RQ2: How well does CTINexus perform in cybersecurity triplet extraction
Demo. Num. F1 Score Precision Recall InputLen
@@
To investigate the impact of prompt example numbers, we evaluated 4 configurations: 1, 2, 3, and 4 examples. Our observations show effectiveness plateau when using 2 or 3 examples, while input ICL prompt size increases significantly with more examples. As shown in Table III, increasing the prompt example number from 1 to 2 improves all metrics, particularly recall. However, with 3 examples, precision and F1-score plateau, and recall drops by 1% . With 4 examples, recall improves from 82 to 84%, but precision drops from 93 to 89%. This contradicts the heuristic that more examples always improve ICL performance but aligns with Chandra et al. [32], noting that each scenario has an optimal number of examples. Additionally, each additional example increases the input length by an average of 603 tokens, slowing inference speed and increasing computational costs. Thus, our implementation uses two examples in the cybersecurity triplet extraction phase, balancing effectiveness and efficiency.
@@
Permutation F1 Score Precision Recall
@@
To analyze the effect of the permutation method for selected examples, we examined three strategies: (1) random selection and sorting (random), (2) selection based on kNN similarity and sorting in ascending order (kNN-ascend), and (3) selection based on kNN similarity and sorting in descending order (kNN-descend). These methods were chosen to explore the impact of recency bias in LLMs, which suggests that models give more weight to examples placed nearer to the query [58]. The random method serves as a baseline, while kNN-ascend and kNN-descend test the influence of example order based on similarity. As shown in Table IV, kNN-ascend outperforms other methods across all metrics, indicating the presence of recency bias and its potential for improving results. Consequently, we adopted kNN-ascend for CTINexus and recommend arranging prompt examples in ascending order of similarity as a universal strategy for other ICL applications.
@@
Backbone F1 Score Precision Recall
@@
The emergence of ICL is closely associated with the substantial parameter counts of LLMs. To assess CTINexus’s generalizability across different backbone models, we evaluate its performance on representative closed-source LLMs, GPT-3.5 and GPT-4, and leading open-source LLMs, Llama3 and Qwen2.5. As shown in Table V, CTINexus achieves over a 10% improvement in both recall and precision when using GPT-4 compared to GPT-3.5-turbo. This underscores the importance of leveraging larger models to fully exploit ICL’s potential within CTINexus’s framework. For Qwen2.5 and Llama3, due to computational resource limitations, we deployed their 72B and 70B parameter versions, respectively. As shown in Table V, both Qwen2.5 and Llama3 demonstrate performance generally comparable to GPT-3.5-turbo. Specifically, Qwen2.5 exhibits a 1.4% higher recall but a 0.9% lower precision compared to Llama3. GPT-4 excelled in both precision and recall among all evaluated backbones. Therefore, all subsequent experiments will employ GPT-4 as the default base model.
5.4 RQ3: How well does CTINexus perform in knowledge graph construction?
@@ 5.4.1 Hierarchical Entity Alignment
As described in Section 4, for entity alignment, we first apply ICL to perform coarse-grained grouping of entities based on their types. We then vectorize these entities into high-dimensional embeddings and conduct fine-grained merging based on their semantic similarity. In the following, we present a series of experiments to investigate the impact of different configurations in entity grouping and entity merging, aiming to identify the optimal combination.
@@
Model Config. Acc Micro-F1 Macro-F1
@@
We assessed the impact of demonstration example numbers on ICL through comparative experiments with four example quantities: 1, 4, 8, and 12. Additionally, we evaluated the performance of two LLMs, GPT-4 and GPT-3.5, across different model sizes. Notably, the performance showed no significant improvement once the number of examples exceeded 12, so these results are excluded from the table. Our evaluation methodology used accuracy, macro-F1, and micro-F1 metrics, consistent with previous text classification studies [63]. The experimental results, shown in Table VI, indicate that GPT-4 consistently outperforms GPT-3.5 across all demonstration number hierarchies. Remarkably, GPT-4 with 1 demonstration yields better results than GPT-3.5 with 12 demonstrations. Both models show substantial improvements when increasing from one to eight demonstrations, but a saturation trend appears when the number of examples exceeds eight. This trend is especially evident in GPT-4, where all three metrics slightly decrease as demonstration numbers increase from eight to twelve.
@@
Threshold F1 Score Precision Recall EntitiyNum
@@
Model F1 Score Precision Recall EntitiyNum
@@
Model Config. F1 Score Precision Recall
@@
TABLE X: Token and time costs of CTINexus across different modules.
Ontology-Aligned Security Triplet Extraction Hierarchical Entity Alignment ICL-Enhanced Long-Distance Relation Prediction Pipeline
Input Output Overall Time Input Output Overall Time Time Input Output Overall Time Token Cost Time
CTINexus w/ GPT-4 0.0246 0.0117 0.0364 11.0905 0.0158 0.0236 0.0393 26.1590 5.9887 0.0644 0.0083 0.0728 24.2483 0.1485 67.4865
CTINexus w/ GPT-3.5 0.0007 0.0006 0.0013 5.9824 0.0008 0.0010 0.0018 10.6606 5.9887 0.0033 0.0005 0.0038 9.5013 0.0069 32.1330
@@
The entity merging module applies a text embedding model to vectorize candidate entities grouped by the entity grouping module and uses a merging threshold to identify equivalent entities. Our evaluation focuses on the selection of the embedding model and the determination of the merging threshold. We use OpenAI’s third-generation embedding models, text-embedding-3-small and text-embedding-3-large, which differ in vector size and represent the latest state-of-the-art general-purpose models. In addition, we also compare with SecureBERT[27], a cybersecurity-specific embedding model based on the RoBERTa architecture pre-trained on a large corpus of cybersecurity data. We consider merging thresholds of 0.4, 0.5, 0.6, and 0.7. Besides common metrics for entity alignment, we introduce Num_ent, which records the number of entities after alignment.
The experimental results are shown in Table VII and Table VIII. Threshold values of 0.4, 0.5, and 0.6 all achieve a 100% recall rate, indicating the algorithm’s ability to detect all entities that should be merged. However, lower thresholds can erroneously merge non-equivalent entities based on the Num_ent and precision metrics. The highest precision is observed when the merging threshold is 0.6. Increasing the threshold to 0.7 maintains precision but significantly reduces recall, suggesting overly fine granularity that misclassifies equivalent entities as distinct. Regarding embedding models, text-embedding-3-large demonstrates the best performance, with text-embedding-3-small showing similar results. SecureBERT, despite its high recall, struggles to correctly cluster entities, as reflected in its low precision and Num_ent scores. This may be due to the smaller size of RoBERTa compared to the text-embedding-3 models, leading to less accurate entity distinction.
5.4.2 ICL-Enhanced Relation Prediction
As mentioned in Section 4, we compose ICL prompts to guide LLMs in inferring relations between disconnected sub-graphs using the provided examples and context. We evaluated different ICL settings by varying the number of demonstration examples (1, 2, and 3) and the sizes of backbone models. Additionally, we examined the effectiveness of zero-shot learning, where the LLM infers relationships of given entities without demonstration examples. Zero-shot learning results are excluded from previous ICL experiments due to poor performance. The better performance in implicit relation inference compared to other tasks in CTINexus could be that relation prediction aligns more closely with general NLP tasks. Unlike triplet extraction or entity alignment, which require domain-specific knowledge in the cybersecurity context, relation prediction relies more on LLMs’ general ability to infer connections between entities based on linguistic cues in the text. This makes relation prediction less dependent on specialized domain knowledge and more aligned with the LLM’s general language understanding capabilities.

### v2

@@ 5 Evaluation
To comprehensively study the performance of CTINexus in various phases of CSKG construction, we set the following research questions:
@@ 1. RQ1:
    How does CTINexus compare with existing methods for CTI knowledge extraction?
@@ 2. RQ2:
    How do different settings in CTINexus affect the cybersecurity triplet extraction?
@@ 3. RQ3:
    How do different settings in CTINexus affect the entity alignment and relation prediction?
@@ 4. RQ4:
    How well does CTINexus perform in end-to-end CSKG construction?
  5. RQ5:
    How well does CTINexus adapt to different CSKG ontologies?
  6. RQ6:
@@
5.1 Dataset and Metrics
Existing datasets primarily benchmark cybersecurity triplet extraction but do not comprehensively support other tasks in our pipeline. Additionally, their CTI reports are often outdated. For instance, LADDER’s dataset [32] includes reports only from 2010 to 2021. To address these limitations, we curated a new dataset specifically for evaluating CTINexus across cybersecurity triplet extraction, hierarchical entity alignment, and long-distance relation prediction tasks. Our dataset consists of 150 recent CTI reports (May 2023 onwards) from 10 reputable sources, such as Trend Micro, Symantec, and The Hacker News.
Annotations
Annotations were performed via a structured four-step approach: (1) Annotating cybersecurity entities and their types. (2) Identifying explicit semantic relations among entities, organized into JSON-formatted triplets. (3) Grouping entities by type and merged mentions referring to identical concepts. (4) Identifying central entities and summarizing implicit inter-entity relations.
To ensure quality and reduce annotation bias, three PhD students with expertise in threat intelligence independently conducted annotations, with the third serving as an arbiter to resolve conflicts. Inter-annotator agreement, measured by Cohen’s kappa [62], yielded scores of 0.80 (triplet extraction), 0.78 (entity alignment), and 0.61 (relation prediction), averaging 0.73, indicating substantial agreement. This rigorous annotation procedure resulted in 3,682 entity mentions, 2,039 unique entities, and 1,982 relationships, enabling comprehensive evaluation of CTINexus’s performance in constructing accurate cybersecurity knowledge graphs
Metrics
For Phase 1 (triplet extraction) and Phase 3 (relation prediction), performance is measured at the triplet level. A predicted <subject, relation, object> counts as a true positive when the three elements semantically match a gold-standard triplet, treating active- and passive-voice variants as equivalent. Predicted triplets with no gold counterpart are false positives, and gold triplets that are not predicted are false negatives. Phase 2 (entity alignment) is assessed at the entity level: a mention is correct if it is (i) assigned the right MALOnt type during coarse-grained grouping, and (ii) merged with the gold mention that refers to the same cybersecurity entity during fine-grained clustering.
@@ 5.1 Dataset and Metrics
5.2 RQ1: How does CTINexus compare against existing CTI knowledge extraction methods?
We evaluated CTINexus against two state-of-the-art baselines, EXTRACTOR [74] and LADDER [32], representing syntactic-analysis-based and fine-tuning-based approaches, respectively.
@@ 5.2 RQ1: How does CTINexus compare against existing CTI knowledge extraction methods?
Several methodological challenges were addressed to enable fair comparison. For EXTRACTOR, we adapted its output to our broader ontology using CTINexus’s coarse-grained entity grouping module. For LADDER, we addressed two key differences: (1) LADDER uses a word-level annotation format, where each token is labeled with its target class. In contrast, our dataset follows an end-to-end report-to-triplet format, where the entire report is input, and the label is a set of extracted triplets. (2) LADDER uses a simplified ontology derived from MALONT, which includes only 10 entity types, a subset of the entity types used in our ontology. To facilitate comparison, we developed scripts to tokenize our data and convert our manually annotated datasets into LADDER’s word-level format. To ensure a fair comparison with LADDER, we merged our training set with LADDER’s in a 5:1 ratio, maintaining their original training/validation split. We also replaced LADDER’s test set with ours to ensure consistent evaluation on the same data. We compare with LADDER solely on named entity extraction performance, as our method focuses on open relation extraction, while LADDER targets relation classification within fixed categories.
Table I demonstrates that CTINexus outperforms EXTRACTOR in all metrics in cybersecurity triplet extraction. The evaluation results in Section 5.3 show that GPT-4 outperforms all other backbone models. Thus, we use GPT-4 as the default backbone model for CTINexus’s implementation.
@@
This superior performance can be attributed to several factors. First, CTINexus leverages the robust context understanding and instruction-following capabilities of LLMs and enhances specificity with kNN-selected demonstration examples for extracting triplets. In contrast, EXTRACTOR employs general fine-tuning to extract semantic roles not specific to any ontology, reducing its accuracy in triplet extraction. Also, the CTI context introduces peculiarities that lead to errors in EXTRACTOR’s semantic role labeling module, which relies on a simple BERT model. For instance, EXTRACTOR might extract a triplet like $\langle$“Androxgh0st malware”, “support”, “numerous functions capable of abusing the Simple Mail Transfer Protocol (SMTP), such as scanning and exploiting exposed credentials and application programming interfaces (APIs), and web shell deployment”$\rangle$, where the object is a long sentence not suitable as a single entity. The object contains multiple entities due to misidentified boundaries. Conversely, CTINexus captures implicit meanings and transforms phrases to be more suitable as entities, resulting in a triplet like $\langle$“Androxgh0st malware”, “supports”, “functions abusing SMTP”$\rangle$.
Table II demonstrates that CTINexus outperforms LADDER in F1-Score, precision, and recall by 26.7%, 17.5%, and 19.5%, respectively. Specifically, LADDER achieved an F1-Score of 71.13%, precision of 78.31%, and recall of 73.94%, which were slightly lower than the numbers reported in LADDER’s original evaluation (75.32%, 79.06%, and 76.98%, respectively). LADDER’s lower performance on our test data compared to its reported values was likely due to a distribution shift. LADDER’s dataset spans 2015 to 2021, while our data is from May 2023 onward. This temporal gap may introduce new patterns, terminologies, or threat vectors that LADDER’s model struggles to generalize to, even when retrained on a mix of old and new data.
@@
The performance disparity between LADDER and CTINexus can be attributed to several factors. First, fine-tuning the model in LADDER may lead to overfitting on the training set. Consequently, when confronted with unseen entities in the test set, the model may struggle to recognize them accurately, potentially misclassifying them or recognizing only parts of the entities. For example, in the sentence “… with a specific focus on WordPress sites”, LADDER extracts only “WordPress” as an application, resulting in ambiguous content. In contrast, CTINexus correctly extracts “WordPress sites”, which more accurately reflects the original context. Second, similar to EXTRACTOR, the LADDER model lacks sufficient contextual understanding. For instance, in the sentence “The victims include Family Day Care Services, a Canadian childcare service”, LADDER incorrectly identifies “Canadian” as a “B-Location”, whereas it should be recognized as a descriptive term for the childcare service. Furthermore, LADDER models relation extraction as relation classification, limited to ten relation classes (i.e., closed-world setting). This constraint restricts the contextual information in the extracted content and hinders the model’s ability to generalize to new CTI data containing different or additional relation classes.
@@
Method F1-Score Precision Recall
@@
Method F1-Score Precision Recall
@@
5.3 RQ2: How do different settings affect the cybersecurity triplet extraction?
To demonstrate the effectiveness of CTINexus in cybersecurity triplet extraction, stemming from the superiority of the ICL paradigm and our specific prompt design, we conducted experiments on different ICL configurations, focusing on four aspects: (1) the number of demonstration examples, (2) the permutation of these examples, (3) the backbone model types, and (4) the prompt formulation and design. By default, CTINexus uses GPT-4 as the model backbone, selects the $k$ most similar prompt examples sorting in ascending order of query similarity.
@@ 5.3 RQ2: How do different settings affect the cybersecurity triplet extraction?
Demo. Num. F1-Score Precision Recall InputLen
@@
To investigate the impact of prompt example numbers, we evaluated 4 configurations: 1, 2, 3, and 4 examples. Our observations show effectiveness plateaus when using 2 or 3 examples, while input ICL prompt size increases significantly with more examples. As shown in Table III, increasing the prompt example number from 1 to 2 improves all metrics, particularly recall. However, with 3 examples, precision and F1-score plateau, and recall drops by 1%. With 4 examples, recall improves from 82 to 84%, but precision drops from 93 to 89%. This contradicts the heuristic that more examples always improve ICL performance but aligns with Chandra et al. [34], noting that the optimal number of examples varies across scenarios. Additionally, each additional example increases the input length by an average of 603 tokens, slowing inference speed and increasing computational costs. Thus, our implementation uses two examples in the cybersecurity triplet extraction phase, balancing effectiveness and efficiency.
@@
Permutation F1-Score Precision Recall
@@
To analyze the effect of the permutation method for selected examples, we examined three strategies: (1) random selection and sorting (random), (2) selection based on kNN similarity and sorting in ascending order (kNN-ascend), and (3) selection based on kNN similarity and sorting in descending order (kNN-descend). These methods were chosen to explore the impact of recency bias in LLMs, which suggests that models give more weight to examples placed nearer to the query [60]. The random method served as a baseline, while kNN-ascend and kNN-descend tested the influence of example order based on similarity. As shown in Table IV, kNN-ascend outperforms other methods across all metrics, indicating the presence of recency bias and its potential for improving results. Consequently, we adopt kNN-ascend for CTINexus and recommend arranging prompt examples in ascending order of similarity as a universal strategy for other ICL applications.
@@
Backbone F1-Score Precision Recall
@@
The emergence of ICL is closely associated with the substantial parameter counts of LLMs. To assess CTINexus’s generalizability across different backbone models, we evaluated its performance on representative closed-source LLMs, GPT-3.5 and GPT-4, and leading open-source LLMs, Llama3 and Qwen2.5. As shown in Table V, CTINexus achieves over a 10% improvement in both recall and precision when using GPT-4 compared to GPT-3.5-turbo. This underscores the importance of leveraging larger models to fully exploit ICL’s potential within CTINexus’s framework. For Qwen2.5 and Llama3, due to computational resource limitations, we deployed their 72B and 70B parameter versions, respectively. As shown in Table V, both Qwen2.5 and Llama3 demonstrate performance generally comparable to GPT-3.5-turbo. Specifically, Qwen2.5 exhibits a 1.4% higher recall but a 0.9% lower precision compared to Llama3. GPT-4 excels in both precision and recall among all evaluated backbones. Therefore, all subsequent experiments employ GPT-4 as the default base model.
Prompt formulation and design
Adhering to our “one-CTI, one-inference” principle, CTINexus demonstrates exceptional data efficiency. We compare two instruction strategies for prompt design: a multi-round QA-based approach (illustrated in Fig. 4 and based on [76]) and our end-to-end prompting approach. In the multi-round method, the extraction prompt comprises three parts: an instruction outlining the extraction task, a context containing the CTI report, and a question requesting specific entities or relations. Consequently, there are $O$ entity extraction prompts and $\frac{N\times(N-1)}{2}$ prompts for relation extraction, where $O$ denotes the number of ontology entity types and $N$ denotes the number of extracted entities. In contrast, CTINexus consolidates these steps into a single inference, extracting all entities and relations in one round with a uniform prompt template. An evaluation of token usage revealed that end-to-end prompting reduces input and output token consumption by 98.9% and 97.3%.
In addition, when designing CTINexus, we iteratively refined the prompt instructions and identified three critical features that significantly boost performance: (1) constraining both output format and content, (2) simulating a role-playing scenario, and (3) placing the instruction text at the beginning of the prompt, before any demonstrations. Compared to a baseline vanilla prompt without these designs, our refined prompt improved F1-score, precision, and recall by 20.57%, 15.28%, and 24.73%, respectively.
5.4 RQ3: How do different settings affect the entity alignment and relation prediction?
@@ 5.4.1 Hierarchical Entity Alignment
As described in Section 4, for entity alignment, we first apply ICL to perform coarse-grained grouping of entities based on their types. We then vectorize entities into high-dimensional embeddings and conduct fine-grained merging based on semantic similarity. In the following, we present a series of experiments to investigate the impact of different configurations in entity grouping and merging, aiming to identify the optimal combination.
@@
Model Config. Acc. Micro-F1 Macro-F1
@@
We assessed the impact of demonstration example numbers on ICL through comparative experiments with four example quantities: 1, 4, 8, and 12. Additionally, we evaluated the performance of two LLMs, GPT-4 and GPT-3.5, across different model sizes. Notably, the performance showed no significant improvement once the number of examples exceeded 12, so these results were excluded from the table. Our evaluation methodology uses accuracy, macro-F1, and micro-F1 metrics, consistent with previous text classification studies [66]. The experimental results, shown in Table VI, indicate that GPT-4 consistently outperformed GPT-3.5 across all demonstration number hierarchies. Remarkably, GPT-4 with 1 demonstration yields better results than GPT-3.5 with 12 demonstrations. Both models show substantial improvements when increasing from one to eight demonstrations, but a saturation trend appears when the number of examples exceeds eight. This trend is especially evident in GPT-4, where all three metrics slightly decrease as demonstration numbers increase from eight to twelve.
@@
Threshold F1-Score Precision Recall EntitiyNum
@@
Model F1-Score Precision Recall EntitiyNum
@@
Model Config. F1-Score Precision Recall
@@
TABLE X: Averaged token cost and time cost per CTI report for each module and the overall CTINexus pipeline, comparing GPT-4 and GPT-3.5.
Backbone Cybersecurity Triplet Extraction Hierarchical Entity Alignment Long-Distance Relation Prediction Overall CTINexus Pipeline
Token Cost (\$/CTI) Time(s/CTI) Token Cost (\$/CTI) Time(s/CTI) Token Cost (\$/CTI) Time(s/CTI) Token Cost (\$/CTI) Time(s/CTI)
GPT-4 0.0364 11.0905 0.0393 26.1590 0.0728 24.2483 0.1485 67.4865
GPT-3.5 0.0013 5.9824 0.0018 10.6606 0.0038 9.5013 0.0069 32.1330
@@
The entity merging module applies a text embedding model to vectorize candidate entities grouped by the entity grouping module and uses a merging threshold to identify equivalent entities. Here, we evaluated different embedding models and merging thresholds. We used OpenAI’s third-generation embedding models, text-embedding-3-small and text-embedding-3-large, which differ in vector size and represent the latest state-of-the-art general-purpose models. In addition, we also compared with SecureBERT[30], a cybersecurity-specific embedding model based on the RoBERTa architecture pre-trained on a large corpus of cybersecurity data. We consider merging thresholds of 0.4, 0.5, 0.6, and 0.7. Besides common metrics for entity alignment, we introduce Num_ent, which records the number of entities after alignment.
The experimental results are shown in Tables VII and VIII. Threshold values of 0.4, 0.5, and 0.6 all achieve a 100% recall rate, indicating the algorithm’s ability to detect all entities that should be merged. However, lower thresholds can erroneously merge non-equivalent entities based on the Num_ent and precision metrics. The highest precision is observed when the merging threshold is 0.6. Increasing the threshold to 0.7 maintains precision but significantly reduces recall, suggesting overly fine granularity that misclassifies equivalent entities as distinct. Regarding embedding models, text-embedding-3-large demonstrates the best performance, with text-embedding-3-small showing similar results. SecureBERT, despite its high recall, struggles to correctly cluster entities, as reflected in low precision and Num_ent scores. This may be due to the smaller size of RoBERTa compared to the text-embedding-3, leading to less accurate entity distinction.
5.4.2 Long-Distance Relation Prediction
As mentioned in Section 4, we compose ICL prompts to guide the LLM in inferring relations between disconnected sub-graphs using demonstration examples and context. We evaluated different ICL settings by varying the number of demonstration examples and the backbone models. Additionally, we examined the effectiveness of zero-shot learning, where the LLM inferred relationships of given entities without demonstration examples. Zero-shot learning results were excluded from previous ICL experiments due to poor performance. The better performance in implicit relation inference compared to other tasks in CTINexus could be that relation prediction aligns more closely with general NLP tasks. Unlike triplet extraction or entity alignment, which require domain-specific knowledge in the cybersecurity context, relation prediction relies more on LLMs’ general ability to infer connections between entities based on linguistic cues. This makes relation prediction less dependent on specialized domain knowledge and more aligned with the LLM’s general language understanding capabilities.